### v1

# On Function-Correcting Codes in the Lee Metric

Abstract
Function-correcting codes are a coding framework designed to minimize redundancy while ensuring that specific functions or computations of encoded data can be reliably recovered, even in the presence of errors. The choice of metric is crucial in designing such codes, as it determines which computations must be protected and how errors are measured and corrected. Previous work by Liu and Liu [6] studied function-correcting codes over $\mathbb{Z}_{2^{l}},\ l\geq 2$ using the homogeneous metric, which coincides with the Lee metric over $\mathbb{Z}_{4}$. In this paper, we extend the study to codes over $\mathbb{Z}_{m},$ for any positive integer $m\geq 2$ under the Lee metric and aim to determine their optimal redundancy. To achieve this, we introduce irregular Lee distance codes and derive upper and lower bounds on the optimal redundancy by characterizing the shortest possible length of such codes. These general bounds are then simplified and applied to specific classes of functions, including Lee-local functions, Lee weight functions, and Lee weight distribution functions, leading to improved some bounds compared to those of Liu and Liu [6] over $\mathbb{Z}_{4}$ and generalize the other bounds over $\mathbb{Z}_{m}$ in the Lee metric.
Keywords: Function-correcting codes, error-correcting codes, Lee metric, redundancy.
Mathematics subject classification: 94B60, 94B65.
1 Introduction
The Lee metric was first introduced by Lee in 1958 as a way to measure distances over cyclic alphabets such as $\mathbb{Z}_{m}$. Later, Golomb and Welch[3], Berlekamp [1], and Chiang [16] developed some of the earliest codes specifically designed for the Lee metric, recognizing its importance in modeling error patterns for non-binary and phase-based communication systems. A method for constructing Lee metric codes over arbitrary alphabet sizes using the elementary concepts of module theory was presented in [12]. These foundational works laid the groundwork for much of the modern research on coding over $\mathbb{Z}_{m}$ using the Lee metric. The Lee metric is often preferred over the Hamming metric in coding theory, particularly when working with codes over rings like $\mathbb{Z}_{m}$ (instead of binary fields), because it better captures the geometry and error characteristics of non-binary alphabet. This motivates the study of function-correcting codes for the Lee metric, as codes optimized for this metric can leverage the modular structure of $\mathbb{Z}_{m}$ to achieve lower redundancy and enhanced error correction performance for the same alphabet size. Several studies have shown that Lee metric based codes can yield tighter redundancy bounds compared to their Hamming metric counterparts, making them particularly effective for applications involving non-binary alphabets and structured, incremental error patterns.
Lenz et al. [5] first introduced the framework of function-correcting codes (FCC) in the Hamming metric, establishing the foundational theory for correcting function evaluations rather than raw symbols. This framework was later extended to other metrics to capture different error models. Specifically, Xia et al. [15] generalized FCC to the symbol-pair metric, while Singh et al. [13] extended it further to the b-symbol metric. Building on these developments, in [14], the authors introduced locally b-symbol codes and investigated the redundancy of the corresponding function-correcting b-symbol codes. Several related advances complement these efforts: Ge et al. [2] studied optimal redundancy bounds for such codes, Ly et al. [7] proposed a conjecture concerning their asymptotic performance, and Premlal et al. [8] developed a function correcting framework for linear functions. In the context of locality, Rajput et al. [9] investigated locally function-correcting codes for the Hamming metric, while Sampath et al. [11] contributed further generalizations of FCC to specialized coding scenarios. Together, these works form a cohesive progression from classical FCC in the Hamming setting to rich variants supporting diverse metrics, locality, and function-specific error corrections.
Function-correcting codes with respect to the homogeneous metric were proposed in [6], where the authors introduced function-correcting codes with homogeneous distance (FCCHDs) as a generalization of function-correcting codes under the Hamming metric. Since both Hamming weight and Lee weight over $\mathbb{Z}_{4}$ can be regarded as special cases of homogeneous weights, the results in [6] naturally extend to the Lee metric over $\mathbb{Z}_{4}$. Motivated by this fact, we investigate the construction and properties of function-correcting codes for the Lee metric over $\mathbb{Z}_{m}$, where $m\geq 2$ is an arbitrary positive integer. We derive upper and lower bounds on the optimal redundancy of function-correcting Lee metric codes for arbitrary functions by employing shortest-length irregular Lee distance codes in Section 3. In Section 4, we provide several bounds on the shortest-length irregular Lee distance codes analogous to the Plotkin and Gilbert-Vashamov bounds. As a result, we improve the bound derived in [6, Lemma 3.2] over $\mathbb{Z}_{4}$. Furthermore, we introduce the concept of locally Lee functions, and using general theoretical results, establish bounds on the optimal redundancy for this class of functions in Section 5. In particular, we demonstrate that the lower bound on optimal redundancy is achievable for locally binary Lee functions. Additionally, we investigate the locality properties of specific functions, the Lee weight function and the Lee weight distribution function, and determine bounds on their optimal redundancy in Section 6. The bounds on the optimal redundancy of function-correcting codes can be obtained as a direct consequence of our general results. In particular, by substituting $m=2$ and $m=4$ into Lemma 4.1, we get the specific cases corresponding to codes in the Hamming distance over $\mathbb{Z}_{2}$ in [5] and the homogeneous distance over $\mathbb{Z}_{4}$ in [6], respectively. This shows that the well-known bounds for these two settings follow naturally from our general framework, without requiring any additional arguments.
2 Preliminaries
In this section, we discuss basic definitions and results that will be used throughout the paper. We denote the residue ring modulo $m$ by $\mathbb{Z}_{m}$, where $m$ is a positive integer greater than or equal to $2$. We consider the elements of $\mathbb{Z}_{m}$ as $10,1,\dots,m-$ 1 unless otherwise specified. For a positive integer $k$, $\mathbb{Z}_{m}^{k}$ denotes the set of all $k$-tuples over $\mathbb{Z}_{m}$.
Definition 2.1.
[4, Lee distance metric] For $u=(u_{1},u_{2},\dots,u_{k}),v=(v_{1},v_{2},\dots,v_{k})\in\mathbb{Z}_{m}^{k}$, the Lee distance between $u$ and $v$ is defined as
$d_{L}(u,v)=\sum_{i=1}^{k}w_{L}(u_{i}-v_{i}),$
where $w_{L}(u_{i}-v_{i})=\min\{u_{i}-v_{i},v_{i}-u_{i}\}\pmod{m}$.
If we denote the elements of $\mathbb{Z}_{m}$ as follows
$\begin{split}-\frac{m-1}{2},-\frac{m-3}{2},\dots,-1,0,1,\dots,\frac{m-1}{2}&\text{ if }m\text{ is odd}\\ -\frac{m-2}{2},\dots,-1,0,1,\dots,\frac{m-2}{2},\frac{m}{2}&\text{ if }m\text{is even},\end{split}$
then the Lee weight of an element $x_{i}\in\mathbb{Z}_{m}$ is equivalently defined as (see [16])
$w_{L}(x_{i})=|x_{i}|.$ (2.1)
The set $\mathbb{Z}_{m}^{k}$ is called the message set. An encoding function $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+r}$ defined as $\mathcal{E}(u)=(u,p(u))$, where $p(u)\in\mathbb{Z}_{m}^{r}$ is the redundancy vector. The set
$C=\{\mathcal{E}(u)\in\mathbb{Z}_{m}^{k+r}\mid u\in\mathbb{Z}_{m}^{k}\}$
is called a code of length $k+r$ over $\mathbb{Z}_{m}$ with redundancy $r$, and $\mathcal{E}(u)$ is called a codeword in $C$. The minimum Lee distance of $C$ is defined as
$d_{L}(C)=\min\{d_{L}(\mathcal{E}(u),\mathcal{E}(v))\mid u,v\in\mathbb{Z}_{m}^{k},u\neq v\}.$
A ball centered at $u\in\mathbb{Z}_{m}^{k}$ of radius $\rho$ is defined as
$B_{L}(u,\rho)=\{v\in\mathbb{Z}_{m}^{k}\mid d_{L}(u,v)\leq\rho\}.$
The cardinality of a ball in $\mathbb{Z}_{m}$ with a given radius is independent of the choice of the center. We denote $V_{L}(r,\rho)=|\{x\in\mathbb{Z}_{m}^{r}:d_{L}(x,c)\leq\rho,\text{for some }c\in\mathbb{Z}_{m}^{r}\}|$, the cardinality of a ball of radius $\rho$ in $\mathbb{Z}_{m}^{r}$. As in the case of the Hamming metric, the sizes of balls appear in the Gilbert-Varshamov bound. Later, we derive a similar bound for function-correcting codes in the Lee metric. The following proposition is therefore useful.
Proposition 2.2.
[10, Proposition 10.10] Let $\rho<m/2$. Then the size of a ball in $\mathbb{Z}_{m}^{r}$ is given by
$V_{L}(r,\rho)=\sum_{j=0}^{r}2^{j}\binom{r}{j}\binom{\rho}{j},$
where $\binom{\rho}{j}=0$ whenever $j>\rho$.
3 Function-correcting Lee metric codes
In error-correcting codes, we add redundancy to each message to enable the receiver to detect and correct any errors that affect the entire message during transmission. In contrast, in function-correcting codes, we need to correct messages with different function values. It allows for more efficient encoding. This approach reduces redundancy when several messages share the same function value. As a result, function-correcting codes offer a targeted and resource-efficient alternative to classical error correction. In this section, we formally define the function-correcting Lee metric code (in short, FCLMC) and establish an equivalence between the function-correcting Lee metric code and irregular-distance codes in the Lee metric. We discuss relations and bounds on the optimal redundancy of an $(f,t)$-FCLMC. Throughout the paper, we assume that a codeword $\mathcal{E}(u)$ is transmitted through a noisy channel, the received word is $a\in\mathbb{Z}_{m}^{k+r}$ with $d_{L}(\mathcal{E}(u),a)\leq t$, that is, at most $t$ errors occur during transmission. We define function-correcting Lee metric codes as follows.
Definition 3.1.
An encoding function $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+r}$ with $\mathcal{E}(u)=(u,p(u))$ defines a function-correcting Lee metric code with redundancy $r$ if $d_{L}(\mathcal{E}(u),\mathcal{E}(v))\geq 2t+1$ holds for all $u,v\in\mathbb{Z}_{m}^{k}$ with $f(u)\neq f(v)$.
Definition 3.2.
The optimal redundancy $r_{f}^{L}(k,t)$ for a given function $f:\mathbb{Z}_{m}^{k}\to Im(f)$ is defined as
$r_{f}^{L}(k,t)=\min\{r\mid\exists(f,t)-\text{FCLMC with redundancy }r\}.$
Any classical error-correcting code with the minimum Lee distance $12t+$ 1, defines an $(f,t)$-FCLMC for any function $f:\mathbb{Z}_{m}^{k}\to Im(f)$. For an integer $n$, we denote $[n]=\{1,2,\dots,n\}$ and $[n]^{+}=\max\{n,0\}$. Define the distance matrix of a given function $f$ as follows.
Definition 3.3.
Let $u_{1},u_{2},\dots,u_{M}\in\mathbb{Z}_{m}^{k}$. Define the distance requirement matrix $D_{f}^{L}(t,u_{1},u_{2},\dots,u_{M})$ of order $N\times N$ in the Lee metric for a given function $f$ as follows:
$[D_{f}^{L}(t,u_{1},u_{2},\dots,u_{M})]_{ij}=\begin{cases}[2t+1-d_{L}(u_{i},u_{j})]^{+}&\text{if }f(u_{i})\neq f(u_{j})\\ 0&\text{ otherwise}\end{cases}$
for all $1\leq i,j\leq M$.
Irregular-distance code in the Lee metric is defined as follows.
Definition 3.4.
Let $D$ be an $M\times M$ square matrix over positive integers. A code $C=\{C_{1},C_{2},\dots,C_{M}\}\subseteq\mathbb{Z}_{m}^{r}$ is a called $D$-code in Lee metric if $d_{L}(C_{i},C_{j})\geq[D]_{ij}$ for all $1\leq i,j\leq M$.
Let $N_{L}(D)$ be the smallest integer $r$ such that there exists a $D$-code of length $r$. When $[D]_{ij}=d$ for all $i\neq j$, we denote $N_{L}(M,d)$. Now, we prove a fundamental relationship between the redundancy of optimal FCLMCs and the irregular distance codes.
Theorem 3.5.
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a function and $\mathbb{Z}_{m}^{k}=\{u_{1},u_{2},\dots,u_{m^{k}}\}$. Then
$r_{f}^{L}(k,t)=N_{L}(D_{f}^{L}(t,u_{1},u_{2},\dots,u_{m^{k}})).$
Proof.
Let $r_{f}^{L}(k,t)=r$ and $\mathcal{E}(u)=(u,p(u))$ be the corresponding encoding function. Let $C=\{p(u_{1}),p(u_{2}),\dots,p(u_{m^{k}})\}\subseteq\mathbb{Z}_{m}^{r}$ be a code of length $r$.
We claim that $C$ is a $D_{f}^{L}(t,u_{1},u_{2},\dots,u_{m^{k}})$-code, that is,
$N_{L}(D_{f}^{L}(t,u_{1},u_{2},\dots,u_{m^{k}}))\leq r.$
If $C$ is not a $D_{f}^{L}(t,u_{1},u_{2},\dots,u_{m^{k}})$-code, then there exist two distinct redundancy vectors $p(u_{i})$ and $p(u_{j})$ with
$0<d_{L}(p(u_{i}),p(u_{j}))<[D_{f}^{L}(t,u_{1},u_{2},\dots,u_{m^{k}})]_{ij}=2t+1-d_{L}(u_{i},u_{j}).$
This implies that $f(u_{i})\neq f(u_{j})$ and
$d_{L}(\mathcal{E}(u_{i}),\mathcal{E}(u_{j}))=d_{L}(u_{i},u_{j})+d_{L}(p(u_{i}),p(u_{j}))<2t+1,$
which is a contradiction to the fact $\mathcal{E}$ defines an $(f,t)$-FCLMC.
Conversely, let $N_{L}(D_{f}^{L}(t,u_{1},u_{2},\dots,u_{m^{k}}))=N$ and $C=\{C_{1},C_{2},\dots,C_{m^{k}}\}$ be a $D_{f}^{L}(t,u_{1},u_{2},\dots,u_{m^{k}})$-code of length $N$. Define an encoding function $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+N}$ defined as $\mathcal{E}(u_{i})=(u_{i},p(u_{i}))$, where $p(u_{i})=C_{i}$ for all $1\leq i\leq m^{k}$. For $f(u_{i})\neq f(u_{j})$, we have
$\begin{split}d_{L}(\mathcal{E}(u_{i}),\mathcal{E}(u_{j}))&=d_{L}(u_{i},u_{j})+d_{L}(C_{i},C_{j})\\ &\geq d_{L}(u_{i},u_{j})+[D_{f}^{L}(t,u_{1},u_{2},\dots,u_{m^{k}}]_{ij}\\ &=d_{L}(u_{i},u_{j})+2t+1-d_{L}(u_{i},u_{j})=2t+1.\end{split}$
Therefore, $\mathcal{E}$ defines an $(f,t)$-FCLMC. Thus $r_{f}^{L}(k,t)\leq N$. This completes the proof. ∎
Corollary 3.6.
Let $\{u_{1},u_{2},\dots u_{M}\}\subseteq\mathbb{Z}_{m}^{k}$ with $M\leq m^{k}$. Then the optimal redundancy of an $(f,t)$-FCLMC satisfies
$r_{f}^{L}(k,t)\geq N_{L}(D_{f}^{L}(t,u_{1},u_{2},\dots,u_{M}))$
for any function $f:\mathbb{Z}_{m}^{k}\to Im(f)$.
Proof.
It is easy to see that
$r_{f}^{L}(k,t)=N_{L}(D_{f}^{L}(t,u_{1},u_{2},\dots,u_{m^{k}}))\geq N_{L}(D_{f}^{L}(t,u_{1},u_{2},\dots,u_{M})),$
since $M\leq m^{k}$. The rest follows from Theorem 3.5. ∎
Next, we define the Lee distance between two distinct function values.
Definition 3.7.
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a function with $Im(f)=\{e_{1},e_{2},\dots,e_{\eta}\}$. Define Lee distance on $Im(f)$ as follows
$d_{L}(e_{i},e_{j})=\min_{u_{1},u_{2}\in\mathbb{Z}_{m}^{k}}d_{L}(u_{1},u_{2})\text{ such that }f(u_{1})=e_{i},f(u_{2})=e_{j}.$
The function-distance matrix of a given function $f$ is defined as follows.
Definition 3.8.
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a function with $Im(f)=\{e_{1},e_{2},\dots,e_{\eta}\}$. Define the function distance matrix with $(ij)$-th entry
$[D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})]_{ij}=\begin{cases}[2t+1-d_{L}(e_{i},e_{j})]^{+}&\text{ if }i\neq j\\ 0&\text{otherwise}.\end{cases}$
One possible approach for constructing $(f,t)$-FCLMCs involves assigning an identical redundancy vector to all message vectors $u$ that correspond to the same function value under a given function. Although this approach is not strictly required for the construction of FCLMCs, it naturally gives an upper bound on the optimal redundancy of FCLMCs as follows.
Theorem 3.9.
For any $f:\mathbb{Z}_{m}^{k}\to Im(f)$ with $Im(f)=\{e_{1},e_{2},\dots,e_{\eta}\}$, we have
$r_{f}^{L}(k,t)\leq N_{L}(D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})).$
Proof.
Let $N(D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta}))=r$, $D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})=D$ and $C=\{C_{i}\mid 1\leq i\leq\eta\}\subseteq\mathbb{Z}_{m}^{r}$ be a $D$-code of length $r$. Define an encoding function $\mathcal{E}(u)=(u,C_{f(u)})$. We claim this encoding defines an $(f,t)$-FCLMC. For this, let $u,v\in\mathbb{Z}_{m}^{k}$ with $e_{i}=f(u)\neq f(v)=e_{j}$. This implies that $C_{f(u)}\neq C_{f(v)}$ and
$\begin{split}d_{L}(\mathcal{E}(u),\mathcal{E}(v))=&d_{L}(u,v)+d_{L}(C_{f(u)},C_{f(v)})\\ \geq&d_{L}(e_{i},e_{j})+[D]_{ij}\\ \geq&d_{L}(e_{i},e_{j})+2t+1-d_{L}(e_{i},e_{j})\\ =&2t+1.\end{split}$
Hence $r_{f}^{L}(k,t)\leq r$. ∎
We have the following corollary from Theorem 3.9 and Corollary 3.6.
Corollary 3.10.
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ with $Im(f)=\{e_{1},e_{2},\dots,e_{\eta}\}$ and a set $u_{1},u_{2},\dots,u_{\eta}$ such that $f(u_{i})=e_{i}$. If $D_{f}^{L}(t,u_{1},u_{2},\dots,u_{\eta})=D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})$, then
$r_{f}^{L}(k,t)=N_{L}(D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})).$
4 Bounds on the shortest length of irregular-distance codes
In this section, we present some bounds on $N_{L}(D)$ that give bounds on the optimal redundancy using results from the previous section. In the next lemma, we give a lower bound on $N_{L}(D)$ for a given Lee distance matrix $D$ which is a generalization of the Plotkin bound for codes with irregular Lee distance requirements.
Lemma 4.1.
Let $D$ be an $M\times M$ Lee distance matrix. Then
$N_{L}(D)\geq\begin{cases}\frac{4}{M^{2}\left\lfloor\frac{m}{2}\right\rfloor}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is even}\\ \frac{4}{(M^{2}-1)\left\lfloor\frac{m}{2}\right\rfloor}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is odd.}\end{cases}$
Proof.
Let $N_{L}(D)=r$. Let $C\subseteq\mathbb{Z}_{m}^{r}$ be a $D$-code of length $r$ and size $M$. Consider an $M\times r$ matrix $A$ whose rows are codewords $\{C_{1},C_{2},\dots,C_{M}\}=C$. First, we derive an upper bound on the sum of the Lee distances between all the codewords.
$\sum_{x,y\in C}d_{L}(x,y)=\sum_{x,y\in C}\sum_{j=1}^{r}d_{L}(x_{j},y_{j})=\sum_{j=1}^{r}\sum_{x,y\in C}d_{L}(x_{j},y_{j}).$
Now, we calculate $\sum_{x,y\in C}d_{L}(x_{j},y_{j}).$ Let $A_{j}$ be the $j^{\mathrm{th}}$-column of $A$. Let $n_{\alpha}$ denote the number of $\alpha\in\mathbb{Z}_{m}$ in $A_{j}$. If $M$ is even, then
$\sum_{x,y\in C}d_{L}(x_{j},y_{j})=\sum_{\alpha,\beta\in\mathbb{Z}_{m}}n_{\alpha}n_{\beta}d_{L}(\alpha,\beta)\leq 2\cdot\frac{M^{2}}{4}\left\lfloor\frac{m}{2}\right\rfloor.$
Maximum attained when $n_{\alpha}=M/2=n_{\beta}$ such that $d_{L}(\alpha,\beta)=\lfloor\frac{m}{2}\rfloor$ for some $\alpha,\beta\in\mathbb{Z}_{m}$ and $n_{\gamma}=0$ for other $\gamma\in\mathbb{Z}_{m}$. Therefore, we have
$\sum_{x,y\in C}d_{L}(x,y)\leq r\frac{M^{2}}{2}\left\lfloor\frac{m}{2}\right\rfloor\ \text{i.e.}\sum_{1\leq i,j\leq N}d_{L}(C_{i},C_{j})\leq r\frac{M^{2}}{2}\left\lfloor\frac{m}{2}\right\rfloor$
Since $C$ is a $D$-code, $d_{L}(C_{i},C_{j})\geq[D]_{ij}$. Thus,
$r\geq\frac{2}{M^{2}\left\lfloor\frac{m}{2}\right\rfloor}\sum_{i,j}[D]_{ij}=\frac{4}{M^{2}\left\lfloor\frac{m}{2}\right\rfloor}\sum_{i,j;i<j}[D]_{ij}.$
When $M$ is odd, then
$\sum_{x,y\in C}d_{L}(x_{j},y_{j})=\sum_{\alpha,\beta\in\mathbb{Z}_{m}}n_{\alpha}n_{\beta}d_{L}(\alpha,\beta)\leq 2\cdot\frac{(M-1)(M+1)}{4}\left\lfloor\frac{m}{2}\right\rfloor.$
The rest proof follows similarly. ∎
In the context of regular-distance codes possessing a minimum distance of $d$, by Lemma 4.1, it follows that
$N(M,d)\geq\frac{2d(M-1)}{M\left\lfloor\frac{m}{2}\right\rfloor}.$
Corollary 4.2.
Let $m=4$. For any Lee distance matrix $D\in\mathbb{N}_{0}^{M\times M}$, we have
$N_{L}(D)\geq\begin{cases}\frac{2}{M^{2}}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is even}\\ \frac{2}{(M^{2}-1)}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is odd.}\end{cases}$
Remark 4.3.
As the homogeneous metric coincides with the Lee metric over $\mathbb{Z}_{4}$. In [6, Lemma 3.2], Liu et al. provided a lower bound on $N_{L}(D)$. The above Corollary 4.2 gives a better lower bound on $N_{L}(D)$ over $\mathbb{Z}_{4}$. Precisely, it improves the bound in [6] by a multiple of $2$ over $\mathbb{Z}_{4}$.
It is well known that the Lee metric coincides with the Hamming metric over $\mathbb{Z}_{2}$. Therefore, [5, Lemma 1] can be derived from Lemma 4.1 by putting $m=2$. We state it in the following corollary.
Corollary 4.4.
Let $m=2$. For any Lee distance matrix $D\in\mathbb{N}_{0}^{M\times M}$, we have
$N_{L}(D)\geq\begin{cases}\frac{4}{M^{2}}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is even}\\ \frac{4}{(M^{2}-1)}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is odd.}\end{cases}$
Analogous to the Gilbert-Varshamov bound, we derive the following lemma for irregular-distance codes in the Lee metric.
Lemma 4.5.
Let $D\in\mathbb{N}^{M\times M}$ be a distance matrix. Then
$N_{L}(D)\leq\min_{r\in\mathbb{N}}\left\{r:m^{r}>\max_{j\in[M]}\sum_{i=1}^{j-1}V_{L}(r,[D]_{ij}-1)\right\},$
where $V_{L}(r,\rho)$ denotes the cardinality of the ball in $\mathbb{Z}_{m}^{r}$ of radius $\rho$.
Proof.
We construct a $D$-code of length $r$. First, select an arbitrary codeword $C_{1}$. For $j=2$, since $m^{r}>V_{L}(r,[D]_{12}-1)$, there exists $C_{2}\in\mathbb{Z}_{m}^{r}$ such that $d_{L}(C_{1},C_{2})\geq[D]_{12}$. For $j=3$, $m^{r}>V_{L}(r,[D]_{13}-1)+V_{L}(r,[D]_{23}-1)$, there exists $C_{3}\in\mathbb{Z}_{m}^{r}$ such that $d_{L}(C_{1},C_{3})\geq[D]_{13}$ and $d_{L}(C_{2},C_{3})\geq[D]_{23}$. Similarly, we chose $C_{4},C_{5},\dots,C_{M}$ such that $d_{L}(C_{i},C_{j})\geq[D]_{ij}$ for $i<j$. Thus, if
$m^{r}>\max_{j\in[M]}\sum_{i=1}^{j-1}V_{L}(r,[D]_{ij}-1),$
then there exist a $D$-code, namely the code $\{C_{1},C_{2},\dots,C_{M}\}$. This completes the proof. ∎
A minor computational error appears in [5, Lemma 4]. We state the corrected version here (see also [15]).
Lemma 4.6.
For any $M,d\in\mathbb{N}$ with $d\geq 10$ and $M\leq d^{2}$, we have
$N_{H}(M,d)\leq\frac{2d-2}{1-2\sqrt{ln(d)/d}},$
where $N_{H}(M,d)$ is the smallest length code such that the code is a $D$-code with respect to the Hamming distance.
Lemma 4.7.
Let $D$ be a distance matrix with $[D]_{ij}=d$ for all $i\neq j$. Then $N_{L}(M,d)\leq N_{H}(M,d)$.
Proof.
The proof follows from the fact that $d_{L}(x,y)\geq d_{H}(x,y)$ for all $x,y\in\mathbb{Z}_{m}$, where $d_{H}(x,y)$ is the Hamming distance between $x$ and $y$. ∎
Using Lemma 4.6 and Lemma 4.7, we have the following.
Corollary 4.8.
For any $M,d\in\mathbb{N}$ with $d\geq 10$ and $M\leq d^{2}$, we have
$N_{L}(M,d)\leq\frac{2d-2}{1-2\sqrt{ln(d)/d}}.$
Let $N_{L}(\eta,d)=\min\{n\mid\exists\text{a code }C\subseteq\mathbb{Z}_{m}^{n}\text{ with }d_{L}(C)\geq 2t,|C|=\eta\}$. That is, $N_{L}(\eta,d)$ is the minimum possible length of a code over $\mathbb{Z}_{m}$ with size $\eta$ and the minimum Lee distance $d$. The following theorem gives lower and upper bounds on the optimal redundancy of FCLMCs.
Theorem 4.9.
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a function with $\mid Im(f)\mid=\eta\geq 2$. Then
$N_{L}(2,2t)\leq r_{f}^{L}(k,t)\leq N_{L}(\eta,2t).$
Proof.
First, we prove that there exists $u,v\in\mathbb{Z}_{m}^{k}$ such that $f(u)\neq f(v)$ and $d_{L}(u,v)=1$. On contrary, suppose that for all $u,v\in\mathbb{Z}_{m}^{k}$ with $d_{L}(u,v)=1$, we have $f(u)=f(v)$. Fix, $u\in\mathbb{Z}_{m}^{k}$, then $f(v)=f(u)$ for all $v\in B_{L}(u,1)$. Next, fix a $u\neq u^{\prime}\in B_{L}(u,1)$, then $f(v)=f(u^{\prime})=f(u)$ for all $v\in B_{L}(u^{\prime},1)$. Similarly, doing same process, we have $f(v)=f(u)$ for all $v\in\mathbb{Z}_{m}^{k}$. This implies $\mid Im(f)\mid=1$, which is a contradiction. Therefore, by Corollary 3.6, we have $r_{f}^{L}(k,t)\geq N_{L}(2,2t)$.
Let $N_{L}(\eta,2t)=N$ and $C=\{C_{i}\mid 1\leq i\leq\eta\}\subseteq\mathbb{Z}_{m}^{N}$ be a code of length $N$ with size $\eta$ and minimum Lee distance $2t$. Let $Im(f)=\{e_{1},e_{2},\dots,e_{\eta}\}$.
Define an encoding function $\mathcal{E}(u)=(u,C_{u})$, where
$C_{u}=\begin{cases}C_{1}&\text{ if }f(u)=e_{1}\\ C_{2}&\text{ if }f(u)=e_{2}\\ \vdots\\ C_{\eta}&\text{ if }f(u)=e_{\eta}.\end{cases}$
We claim that this encoding defines an $(f,t)$-FCLMC. For this, let $u,v\in\mathbb{Z}_{m}^{k}$ with $f(u)\neq f(v)$. Then $C_{u}\neq C_{v}$ and thus,
$\begin{split}d_{L}(\mathcal{E}(u),\mathcal{E}(v))=d_{L}(u,v)+d_{L}(C_{u},C_{v})\geq 1+2t.\end{split}$
Hence $r_{f}^{L}(k,t)\leq N_{L}(\eta,2t)$. ∎
5 FCLMCs for the locally functions
In applications, we utilize specific functions or classes of functions that exhibit desirable properties, enhancing the efficiency of function-correcting codes. These structured functions often simplify encoding and decoding processes. In this section, we investigate one such class, known as locally defined functions. We explore the optimal redundancy of FCLMCs for these functions. This focused study helps in constructing more practical and effective FCLMCs. First, we provide values of $\lambda$ and $\rho$ for which a given function behaves like a locally $(\lambda,\rho)_{L}$-function, and then provide bounds on the optimal redundancy. Moreover, we describe locality and the optimal redundancy of $(f,t)$-FCLMCs for the Lee weight function and the Lee weight distribution function. We start by defining the function ball and locally functions in the Lee metric.
Definition 5.1.
Let $u\in\mathbb{Z}_{m}^{k}$. A function ball centered at $u$ of radius $\rho$ is defined as
$B_{L}^{f}(u,\rho)=\{f(v)\mid v\in B_{L}(u,\rho)\}.$
Definition 5.2.
A function $f:\mathbb{Z}_{m}^{k}\to Im(f)$ is said to be a locally $(\lambda,\rho)_{L}$-function if $\mid B_{L}^{f}(u,\rho)\mid\leq\lambda$ for all $u\in\mathbb{Z}_{m}^{k}$.
When $\lambda=2$, we say the function is a locally binary Lee function. As the optimal redundancy can be obtained by analyzing those message words for which function values are different. The following lemma distinguishes the message words inside a function ball. The proof follows a similar approach to that of the Hamming distance [9]. Therefore, we omit the proof.
Lemma 5.3.
Let $f:\mathbb{F}_{2}^{k}\to Im(f)$ be a locally $(\lambda,\rho)_{L}$-function. Then there exists a map $\tau_{f}^{L}:\mathbb{F}_{2}^{k}\to[\lambda]$ such that $Im(\tau_{f}^{L})\leq[\lambda]$ and for all $u,v\in\mathbb{F}_{2}^{k}$, $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$ if either $u\in B_{L}^{f}(v,\rho)$ or $v\in B_{L}^{f}(u,\rho)$.
Using the above lemma, we construct an $(f,t)$-FCLMC for locally $(\lambda,2t)_{L}$-functions. This will give an upper bound on the optimal redundancy in terms of the existence of classical codes in the Lee metric with prescribed parameters depending on $\lambda$ and $t$.
Theorem 5.4.
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a locally $(\lambda,2t)_{L}$-function. Then, an upper bound on the optimal redundancy of an $(f,t)$-FCLMC is given by
$r_{f}^{L}(k,t)\leq N_{L}(\lambda,2t).$
Proof.
By Lemma 5.3, there exists a map $\tau_{f}^{L}:\mathbb{Z}_{m}^{k}\to[\lambda]$ such that for $u,v\in\mathbb{F}_{2}^{k}$, $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$ whenever $f(u)\in B_{L}^{f}(v,2t)$ or $f(u)\in B_{L}^{f}(u,2t)$. Let $C=\{C_{1},C_{2},\dots,C_{\lambda}\}$ be a code of length $N_{L}(\lambda,2t)$ with $\lambda$ codewords and minimum Lee distance $2t$. Define an encoding $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+N_{L}(\lambda,2t)}$ given by $\mathcal{E}(u)=(u,C_{\tau_{f}^{L}(u)})$. We claim that this encoding defines an $(f,t)$-FCLMC. Let $u,v\in\mathbb{Z}_{m}^{k}$ with $f(u)\neq f(v)$. Then
Case 1: If $d_{L}(u,v)\geq 2t+1$, then
$d_{L}(\mathcal{E}(u),\mathcal{E}(v))\geq d_{L}(u,v)\geq 2t+1.$
Case 2: If $d_{L}(u,v)\leq 2t$, then $f(v)\in B_{L}^{f}(u,2t)$. We have $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$. Thus $C_{\tau_{f}^{L}(u)}\neq C_{\tau_{f}^{L}(v)}$. We have,
$\begin{split}d_{L}(\mathcal{E}(u),\mathcal{E}(v))=d_{L}(u,v)+d_{L}(C_{\tau_{f}^{L}(u)},C_{\tau_{f}^{L}(v)})\geq 2t+1.\end{split}$
Hence $r_{f}^{L}(k,t)\leq N_{L}(\lambda,2t)$. ∎
The following corollary shows that the optimal redundancy of FCLMCs is achieved for locally binary Lee functions.
Corollary 5.5.
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a locally $(2,2t)_{L}$-function. Then $r_{f}^{L}(k,t)=N_{L}(2,2t)$.
Proof.
The Proof follows using Theorem 5.4 and Theorem 4.9. ∎
The above corollary suggests that determining the value $N_{L}(2,2t)$ will give the optimal redundancy of $(f,t)$-FCLMC for locally $(2,2t)_{L}$-functions.
Proposition 5.6.
For positive integers $t$ and $m\geq 2$, we have
$N_{L}(2,2t)=\left\lceil\frac{2t}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil.$
Proof.
Denote $\left\lceil\frac{2t}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil=s$. Consider $c=(0,0,,\dots,0)$ and $d=(\lfloor\frac{m}{2}\rfloor,\lfloor\frac{m}{2}\rfloor,\dots,\lfloor\frac{m}{2}\rfloor)$ of length $s$. Then
$d_{L}(c,d)=\sum_{i=1}^{s}w_{L}\left(0-\left\lfloor\frac{m}{2}\right\rfloor\right)=s\cdot\left\lfloor\frac{m}{2}\right\rfloor\geq 2t.$
Thus, $N_{L}(2,2t)\leq s$.
Conversely, let $c=(c_{1},c_{2},\dots,c_{s-1})$ and $d=(d_{1},d_{2},\dots,d_{s-1})$ be two codewords of length $s-1$. Then, as the maximum Lee weight of an element in $\mathbb{Z}_{m}$ can be $\left\lfloor\frac{m}{2}\right\rfloor$, we have
$d_{L}(c,d)=\sum_{i=1}^{s-1}w_{L}(c_{i}-d_{i})\leq(s-1)\cdot\left\lfloor\frac{m}{2}\right\rfloor<2t.$
Therefore, $N_{L}(2,2t)\geq s$. This completes the proof. ∎
Remark 5.7.
By Proposition 5.6, if $m=2$, then $N_{L}(2,2t)=2t$ and if $m=4$, then $N_{L}(2,2t)=t$. Thus, Proposition 5.6 is consistent with the results obtained in [5] and [6] for the Hamming distance and the homogeneous distance, respectively.
6 FCLMCs for the Lee weight and Lee weight distribution functions
In this section, we explore locality and optimal redundancy of FCLMCs for the Lee weight and the Lee weight distribution functions.
Let $f(u)=w_{L}(u)$ for all $u\in\mathbb{Z}_{m}^{k}$ be Lee wight function. Recall that
$B_{L}(u,\rho)=\{v\in\mathbb{Z}_{m}^{k}\mid d_{L}(u,v)\leq\rho\}.$
Let $v\in B_{L}(u,\rho)$. Then by the triangle inequality, we have
$\begin{split}w_{L}(v)=d_{L}(v,0)\leq d_{L}(v,u)+d_{L}(u,0)\leq\rho+w_{L}(u)\end{split}$
and
$\begin{split}w_{L}(u)=d_{L}(u,0)\leq d_{L}(u,v)+d_{L}(v,0)\leq\rho+w_{L}(v).\end{split}$
Therefore, for any $v\in B_{L}(u,\rho)$, we have
$w_{L}(u)-\rho\leq w_{L}(v)\leq w_{L}(u)+\rho.$ (6.1)
This observation gives the following.
Theorem 6.1.
Let $\rho$ be a positive integer. Then the Lee weight function is a locally $(2\rho+1,\rho)_{L}$-function.
Proof.
Let $f(u)=w_{L}(u)$. By Equation 6.1, $\mid B_{L}^{f}(u,\rho)\mid\leq 2\rho+1$ for all $u\in\mathbb{Z}_{m}^{k}$. ∎
Corollary 6.2.
The optimal redundancy of an $(f,t)$-FCLMC for the Lee weight function $f(u)=w_{L}(u)$ satisfies $r_{f}^{L}(k,t)\leq N_{L}(4t+1,2t)$.
Proof.
By Theorem 6.1, the Lee weight function $w_{L}(u)$ is a locally $(4t+1,2t)_{L}$-function. The rest follows from Theorem 5.4. ∎
We know that the maximum possible Lee weight of an element in $\mathbb{Z}_{m}$ can be $\left\lfloor\frac{m}{2}\right\rfloor$. In fact, for any $u\in\mathbb{Z}_{m}^{k}$, the Lee weight of $u$, $w_{L}(u)\in\{0,1,\dots,k\left\lfloor\frac{m}{2}\right\rfloor\}=Im(w_{L})$. It is easy to see that for any $i,j\in\{0,1,2,\dots,k\}=Im(w_{H})$, there exist $u_{i}=(1^{i},0^{k-i}),u_{j}=(1^{j},0^{k-j})\in\mathbb{Z}_{m}^{k}$ such that $w_{H}(u_{i})=i$, $w_{H}(u_{j})=j$ and $d_{H}(u_{i},u_{j})=|i-j|$. In the Lee metric, finding such elements is non-trivial. In the following lemma, we construct such elements for the Lee metric.
Lemma 6.3.
For all $i\in\{0,1,\dots,k\left\lfloor\frac{m}{2}\right\rfloor\}=Im(w_{L})$, there exist an $u_{i}\in\mathbb{Z}_{m}^{k}$ such that $w_{L}(u_{i})=i$ and $d_{L}(u_{i},u_{j})=|i-j|$.
Proof.
Define $\mathcal{CL}_{0}=\{0\}$ and for $1\leq a\leq k$
$\mathcal{CL}_{a}=\left\{i\in\mathbb{N}\mid(a-1)\left\lfloor\frac{m}{2}\right\rfloor<i\leq a\left\lfloor\frac{m}{2}\right\rfloor\right\}.$
Observe that
$\bigcup_{0\leq a\leq k}\mathcal{CL}_{a}=Im(w_{L}).$
Also, let $u_{0}=(0,0,\dots,0)\in\mathbb{Z}_{m}^{k}$. For $1\leq i\leq k\left\lfloor\frac{m}{2}\right\rfloor$, if $i\in\mathcal{CL}_{a}$, that is, $(a-1)\left\lfloor\frac{m}{2}\right\rfloor<i\leq a\left\lfloor\frac{m}{2}\right\rfloor$, then
$u_{i}=\left(\left\lfloor\frac{m}{2}\right\rfloor,\left\lfloor\frac{m}{2}\right\rfloor,\dots,\left\lfloor\frac{m}{2}\right\rfloor,i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor(a^{\mathrm{th}}\text{-place}),0,0,\dots,0\right)\in\mathbb{Z}_{m}^{k}.$
Note that
$\left(m-\left(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)-(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)=$ $m-i+(a-1)\left\lfloor\frac{m}{2}\right\rfloor-i+(a-1)\left\lfloor\frac{m}{2}\right\rfloor$
$=$ $m+2(a-1)\left\lfloor\frac{m}{2}\right\rfloor-2i$
$\geq$ $m+2(a-1)\left\lfloor\frac{m}{2}\right\rfloor-2a\left\lfloor\frac{m}{2}\right\rfloor$
$=$ $m-\left\lfloor\frac{m}{2}\right\rfloor\geq 0.$
That is,
$m-\left(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)\geq i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor>0.$
Therefore,
$w_{L}\left(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)=\min\left\{i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor,m-\left(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)\right\}=i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor.$
Consequently,
$w_{L}(u_{i})=(a-1)\left\lfloor\frac{m}{2}\right\rfloor+i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor=i.$
Now, we show that $d_{L}(u_{i},u_{j})=|i-j|$. For this,
Case 1: Let $i,j\in\mathcal{CL}_{a}$ for some $a$. Without loss of generality, assume that $j>i$. Then
[…]
$=$ $w_{L}(i-j).$
Observe that $-\left\lfloor\frac{m}{2}\right\rfloor\leq i-j\leq\left\lfloor\frac{m}{2}\right\rfloor$, thus by Eq.2.1, $w_{L}(i-j)=|i-j|$. Hence $d_{L}(u_{i},u_{j})=|i-j|.$
Case 2: Let $i\in\mathcal{CL}_{a}$ and $j\in\mathcal{CL}_{b}$ for some $a\neq b$. Without loss of generality, let $j>i$, then $b>a$ and
$u_{i}-u_{j}=$
[…]
Thus
$d_{L}(u_{i},u_{j})=$ $w_{L}(u_{i}-u_{j})=w_{L}\left(i-a\left\lfloor\frac{m}{2}\right\rfloor\right)+(b-a-1)w_{L}\left(-\left\lfloor\frac{m}{2}\right\rfloor\right)+w_{L}\left(-j+(b-1)\left\lfloor\frac{m}{2}\right\rfloor\right).$
Note that $(a-1)\left\lfloor\frac{m}{2}\right\rfloor)<i\leq a\left\lfloor\frac{m}{2}\right\rfloor$. This implies that $-\left\lfloor\frac{m}{2}\right\rfloor\leq i-a\left\lfloor\frac{m}{2}\right\rfloor\leq 0\leq\left\lfloor\frac{m}{2}\right\rfloor$. Therefore, again by Eq. 2.1,
$w_{L}(i-a\left\lfloor\frac{m}{2}\right\rfloor)=\left|i-a\left\lfloor\frac{m}{2}\right\rfloor\right|=-i+a\left\lfloor\frac{m}{2}\right\rfloor$
$d_{L}(u_{i},u_{j})=$ $w_{L}(u_{i}-u_{j})=w_{L}\left(i-a\left\lfloor\frac{m}{2}\right\rfloor\right)+(b-a-1)w_{L}\left(\left\lfloor\frac{m}{2}\right\rfloor\right)+w_{L}\left(j-(b-1)\left\lfloor\frac{m}{2}\right\rfloor\right)$
$=$ $-i+a\left\lfloor\frac{m}{2}\right\rfloor+b\left\lfloor\frac{m}{2}\right\rfloor-a\left\lfloor\frac{m}{2}\right\rfloor-\left\lfloor\frac{m}{2}\right\rfloor+j-b\left\lfloor\frac{m}{2}\right\rfloor+\left\lfloor\frac{m}{2}\right\rfloor$
$=$ $-i+j=|i-j|\ \ \ \text{ (since }i-j<0,|i-j|=-(i-j)=-i+j).$
This completes the proof. ∎
Using the above lemma, we have the following.
Theorem 6.4.
Let $f(u)=w_{L}(u)$ be the Lee weight function with $Im(w_{L})=\{0,1,\dots,k\left\lfloor\frac{m}{2}\right\rfloor=\eta\}$. Let $D_{w_{L}}(t)$ be an $(\eta+1)\times(\eta+1)$ matrix with $(i,j)$-th entry
$[D_{w_{L}}(t)]_{ij}=\begin{cases}0&\text{ if }i=j\\ [2t+1-|i-j|]^{+}&\text{ if }i\neq j.\end{cases}$
Then $r_{f}^{L}(k,t)=N_{L}(D_{w_{L}}(t))$.
Proof.
Let $e_{i}=i$ for all $i\in Im(w_{L})$. From Definition 3.7, we have
$d_{L}(e_{i},e_{j})=\min_{u_{1},u_{2}\in\mathbb{Z}_{m}^{k}}d_{L}(u_{1},u_{2})\text{ such that }f(u_{1})=e_{i},f(u_{2})=e_{j}.$
Using the triangle inequality, $d_{L}(e_{i},e_{j})\geq|i-j|$. Also, by Lemma 6.3, there exist $u_{1},u_{2}\in\mathbb{Z}_{m}^{k}$, such that $w_{L}(u_{1})=i$, $w_{L}(u_{2})=j$ and $d_{L}(u_{i},u_{j})=|i-j|$. Hence $d_{L}(e_{i},e_{j})=|i-j|$. By Theorem 3.9, we have
$r_{f}^{L}(k,t)\leq N_{L}(D_{w_{L}}(t)).$
Conversely, for all $i\in\left\{0,1,\dots,k\left\lfloor\frac{m}{2}\right\rfloor\right\}$, by Lemma 6.3, there exits $u_{i}\in\mathbb{Z}_{m}^{k}$ with $w_{L}(u_{i})=i$ such that $d_{L}(u_{i},u_{j})=|i-j|$. Then by Corollary 3.6, $r_{f}^{L}(k,t)\geq N_{L}(D_{w_{L}}(t))$. ∎
Example 6.5.
Let $m=4$, $t=2$ and $k=4$. Then $\eta=k\left\lfloor\frac{m}{2}\right\rfloor=8$ and $D_{w_{L}}(2)$ is a $9\times 9$ matrix given by
$\begin{bmatrix}0&4&3&2&1&0&0&0&0\\ 4&0&4&3&2&1&0&0&0\\ 3&4&0&4&3&2&1&0&0\\ 2&3&4&0&4&3&2&1&0\\ 1&2&3&4&0&4&3&2&1\\ 0&1&2&3&4&0&4&3&2\\ 0&0&1&2&3&4&0&4&3\\ 0&0&0&1&2&3&4&0&4\\ 0&0&0&0&1&2&3&4&0\end{bmatrix}.$
Using Theorem 6.4 and Lemma 4.1, we get a lower bound on the optimal redundancy of FCLMCs for the Lee weight function.
Corollary 6.6.
Let $f(u)=w_{L}(u)$ and $\eta=k\left\lfloor\frac{m}{2}\right\rfloor$. For any $k>\left\lceil\frac{t+1}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil$,
$r_{f}^{L}(k,t)\geq\frac{10t^{3}+30t^{2}+20t}{3(t+2)^{2}\left\lfloor\frac{m}{2}\right\rfloor}.$
Proof.
Let $\{C_{1},C_{2},\dots,C_{\eta+1}\}$ be a $D_{w_{L}}(t)$-code. Consider first $t+2$ codewords $C_{1},C_{2},\dots,C_{t+2}$. By Lemma 4.1 and Theorem 6.4,
$r_{f}^{L}(k,t)=$ $N_{L}(D_{w_{L}}(t))\geq\frac{4}{(t+2)^{2}\left\lfloor\frac{m}{2}\right\rfloor}\sum_{1\leq i,j\leq t+2;i<j}[D_{w_{L}}(t)]_{ij}$
$=$ $\frac{4}{(t+2)^{2}\left\lfloor\frac{m}{2}\right\rfloor}\sum_{i=1}^{t+2}\sum_{j=i+1}^{t+2}[D_{w_{L}}(t)]_{ij}$
$=$ $\frac{4}{(t+2)^{2}\left\lfloor\frac{m}{2}\right\rfloor}\left(\sum_{i=0}^{t}(2t-i)(t+1-i)\right)$
$=$ $\frac{4}{(t+2)^{2}\left\lfloor\frac{m}{2}\right\rfloor}\frac{5t^{3}+15t^{2}+10t}{6}$
$=$ $\frac{10t^{3}+30t^{2}+20t}{3(t+2)^{2}\left\lfloor\frac{m}{2}\right\rfloor}.$
Remark 6.7.
Let $f:\mathbb{Z}_{4}^{k}\to Im(f)$ be the Lee weight function. Then by Corollary 6.6, for any $k>\left\lceil\frac{t+1}{2}\right\rceil$,
$r_{f}^{L}(k,t)\geq\frac{5t^{3}+15t^{2}+10t}{3(t+2)^{2}}.$
In [6, Corollary 4.2], the authors proved the same bound for the homogeneous metric. As the homogeneous metric and the Lee metric are the same over $\mathbb{Z}_{4}$. However, Corollary 6.6 gives the bound over $\mathbb{Z}_{m}$ for any positive integer $m$ with respect to the Lee metric.
Now, we describe the locality of the Lee weight distribution function. The Lee weight distribution function with a threshold $T>0$ is defined as
$W_{L}^{T}(u)=\left\lfloor\frac{w_{L}(u)}{T}\right\rfloor.$
The following theorem gives a value of $\lambda$ for which the Lee weight distribution function $W_{L}^{T}$ is a locally $(\lambda,\rho)_{L}$-function.
Theorem 6.8.
Let $\rho$ be a positive integer. Then the Lee weight distribution function $W_{L}^{T}$ is a locally $\left(\left\lfloor\frac{2\rho}{T}\right\rfloor+2,\rho\right)_{L}$-function.
Proof.
Let $f(u)=W_{L}^{T}(u)$. Let $u\in\mathbb{Z}_{m}^{k}$, by Equation 6.1, for any $v\in B_{L}(u,\rho)$, we have
$\left\lfloor\frac{w_{L}(u)-\rho}{T}\right\rfloor\leq f(v)=\left\lfloor\frac{w_{L}(v)}{T}\right\rfloor\leq\left\lfloor\frac{w_{L}(u)+\rho}{T}\right\rfloor.$
Therefore,
$\mid B_{L}^{f}(u,\rho)\mid\leq\left\lfloor\frac{w_{L}(u)+\rho}{T}\right\rfloor-\left\lfloor\frac{w_{L}(u)-\rho}{T}\right\rfloor+1\leq\left\lfloor\frac{(w_{L}(u)+\rho)-(w_{L}(u)-\rho)}{T}\right\rfloor+2,$
that is, $|B_{L}^{f}(u,\rho)|\leq\left\lfloor\frac{2\rho}{T}\right\rfloor+2$. Since $u$ is arbitrary, therefore, $W_{L}^{T}$ is a locally $\left(\left\lfloor\frac{2\rho}{T}\right\rfloor+2,\rho\right)_{L}$-function. ∎
Corollary 6.9.
The optimal redundancy of an $(f,t)$-FCLMC for the Lee weight distribution function $f(u)=W_{L}^{T}(u)$ satisfies $r_{f}^{L}(k,t)\leq N_{L}\left(\left\lfloor\frac{4t}{T}\right\rfloor+2,2t\right)$.
Proof.
By Theorem 6.8, the Lee weight distribution function $W_{L}^{T}(u)$ is a locally $\left(\left\lfloor\frac{4t}{T}\right\rfloor+2,2t\right)_{L}$-function. The rest follows from Theorem 5.4. ∎
7 Conclusion
In this paper, we extended the study of function-correction codes for the Lee metric by introducing function-correcting Lee metric codes (FCLMCs) over the residue ring $\mathbb{Z}_{m}$. FCLMCs require less redundancy compared to classical Lee metric codes. In this encoding, we only need to correct those codewords whose associated function values differ, rather than correcting all erroneous codewords. We derived upper and lower bounds on the optimal redundancy of FCLMCs for general functions utilizing the smallest length irregular Lee distance codes. Moreover, we defined locally Lee functions, and using general results, we obtained bounds on the optimal redundancy for locally Lee functions. We also showed that the lower bound on optimal redundancy can be attained in the case of locally binary Lee functions. We further explored the locality of specific functions, namely the Lee weight function and the Lee weight distribution function, and established bounds on the optimal redundancy for these functions. Our results generalize the corresponding bounds over $\mathbb{Z}_{2}$ given in [5] and over $\mathbb{Z}_{4}$ given in [6]. Moreover, Corollary 4.2 improves the bound obtained in [6] over $\mathbb{Z}_{4}$. Potential directions for future work include extending the analysis of FCLMCs to a specific class of functions of interest and examining the tightness of the bounds obtained in this paper.